1. Solve y = 5e3x + sin x (dy/dx)
  2. The velocity of a body as a function of time is given as ν (t) = 5-2t  + 4, where t is in seconds, and v is in m/s. Solve the acceleration in m/s2 at t = 0.6
1
Expert's answer
2021-10-28T07:01:34-0400

1.


"\dfrac{dy}{dx}=\dfrac{d}{dx}(5e^{3x}+\sin x)=15e^{3x}+\cos x"

2.


"a(t)=\dfrac{dv}{dt}=\dfrac{d}{dt}(5e^{-2t}+4)=-10e^{-2t}"

"a(0.6)=-10e^{-2(0.6)}\ m/s^2"

"=-10e^{-1.2}\ m/s^2\approx-3.012\ m/s^2"